OpenStudy (fools101):

If x = i Root of 6 is one zero of the function: y = x3 + 3x2 + 6x + 18, find the other zeros.?

OpenStudy (fools101):

please help!

OpenStudy (zarkon):

if \[i\sqrt{6}\] is a root then so is \[-i\sqrt{6}\] and your polynomial has the two factors \[(x-i\sqrt{6})(x+i\sqrt{6})\]

OpenStudy (zarkon):

\[(x-i\sqrt{6})(x+i\sqrt{6})=x^2+6\] so \[x^3 + 3x^2 + 6x + 18=(x-a)(x^2+6)\] thus a=-3 ( we need -a*6=18) so the roots are \[-3,i\sqrt{6},-i\sqrt{6}\]

OpenStudy (fools101):

THANK UUUUU!!! ZARKON!!